OpenStudy (anonymous):

factorise and simplify 3(6-2a)+a-3/9-3a

OpenStudy (anonymous):

is anyone able to help with this one?

OpenStudy (anonymous):

[3(6-2a)+a-3] is this all a numerator?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

are sure about all signs of numerator and denomenator?

OpenStudy (anonymous):

numerators are 3(6-2a)+a-3 denomenators are 9-3a

OpenStudy (anonymous):

= ( 18-6a+a-3)/(9-3a) = (15-5a)/(9-3a) = 5( 3-a)/3(3-a) = 5/3

OpenStudy (anonymous):

expand the numerator then take a common factor of 5 in numerator and a common factor of 4 denomenator, you cancel to ger 5/3 , hope that helps.

OpenStudy (anonymous):

thats great, thanks so much

OpenStudy (anonymous):

you are welcome.